OpenStudy (anonymous):

Will fan and medal. What is the value of x? x+15=8

OpenStudy (immanuelv):

x = 8 - 15

OpenStudy (anonymous):

Thank you!

OpenStudy (immanuelv):

you're welcome :D

OpenStudy (anonymous):

@immanuelv the answer is -7 right?

OpenStudy (immanuelv):

you got it! yea its -7

OpenStudy (anonymous):

Alright, thanks again. I may have more questions for you soon. I'm about 20 lessons behind in math, because I dont pay attention haha.

OpenStudy (anonymous):

What is the value of x? 14-x=21 so i'd do 14-21=x? @immanuelv

OpenStudy (immanuelv):

yes!

OpenStudy (anonymous):

Alright, thanks!

OpenStudy (immanuelv):

you're welcome